OpenStudy (anonymous):

\[tan^2 \theta +100 = ?\]? in terms of \[sec\theta\]

OpenStudy (anonymous):

well i guess thee question is wrong :P

OpenStudy (anonymous):

See the Identity is: \[1 + \tan^2 \theta = \sec^2 \theta\]

OpenStudy (anonymous):

so what if its 100 in stead of 1?

OpenStudy (anonymous):

Multiply 10 both sides of the formula that I have written..

OpenStudy (anonymous):

i see...that would make \[10 + 10\tan^2 \theta = 10\sec^2 \theta\] so where from here? because I guess I don't know what \[100 + \tan^2 \theta = \] in terms of sec

OpenStudy (king):

nw u have 2 equations for 10sec^2theta,so over solve...

OpenStudy (anonymous):

No no go slowly.. See original and formula equation have now become: \[\tan^2 \theta + 100 = 10\sec^2 \theta\] --------1 \[10\tan^2 \theta + 10 = 10 \sec^2 \theta\] ----2 Now RHS of both the equations are equal so equate LHS.. Getting?

OpenStudy (anonymous):

shoot me

OpenStudy (anonymous):

i 'm not seeing it....what's wrong with me?

OpenStudy (anonymous):

Okay where is the gun??

OpenStudy (anonymous):

LOOOOLzzzz!!!!!

OpenStudy (anonymous):

Just go with more simple method.. See the formula I have written above.. You know from there the value of sec^2 theta...

OpenStudy (anonymous):

you mean substitute one equation into the other and solve for sec^2theta?

OpenStudy (anonymous):

See the formula is: \[1 + \tan^2 \theta = \sec^2 \theta\] This value of sec^2 theta just put in the question see what you get?

OpenStudy (anonymous):

hold on a sec

OpenStudy (anonymous):

Take your time..

OpenStudy (anonymous):

Have you changed the question??

OpenStudy (anonymous):

Well it's the same question....I just don't know what the RHS equals

OpenStudy (anonymous):

But there you have written 10sec^2 theta earlier??

OpenStudy (anonymous):

10 sec^2theta was my guess

OpenStudy (anonymous):

I followed it with a question mark

OpenStudy (anonymous):

LOL! Sorry I guess I'm confusing the both of us

OpenStudy (anonymous):

So, the original question is: \[\tan^2 \theta + 100 = ???\] Is it??

OpenStudy (anonymous):

yes....do you want me to write the whole question....this is just a small part of it

OpenStudy (turingtest):

definitely need the whole question here

OpenStudy (anonymous):

Yes if you want to then write the whole.. Firstly here subtract 100 from both the sides and tell me what you got??

OpenStudy (anonymous):

\[\int \frac{x^3}{\sqrt(x^2 +100)} dx\]

OpenStudy (anonymous):

LOL no....what does it equal in terms of sec?

OpenStudy (turingtest):

you need to remember the constants in your substitutions! you forgot yours

OpenStudy (anonymous):

x=tan theta dx=sec^2 theta d theta

OpenStudy (anonymous):

i was trying to change the denomitor in terms of sec to get rid of the sqrt

OpenStudy (turingtest):

yeah, but what about the 100? we need to take care of that the sub you want for integrals of the form\[\sqrt{ax^2+b}\]is\[x=\sqrt\frac ba\tan\theta\]look at what happens when you square that and it should make more sense...

OpenStudy (anonymous):

hold on turing

OpenStudy (anonymous):

ok go ahead...i just don't see how we went to \[x=\sqrt\frac ba\tan\theta\]

OpenStudy (turingtest):

because it always works, watch:\[\sqrt{ax^2+b}\]\[x=\sqrt\frac ba\tan\theta\implies x^2=\frac ba\tan^2\theta\]so we get\[\sqrt{a\left(\frac ba\tan^2\theta\right)+b}=\sqrt{b\tan^2\theta+b}=\sqrt b\cdot\sqrt{\tan^2+1}=\sqrt b\sec\theta \]

OpenStudy (anonymous):

I can follow this to problem: \[\sqrt{a\left(\frac ba\tan^2\theta\right)+b}=\sqrt{b\tan^2\theta+b}=\sqrt b\cdot\sqrt{\tan^2+1}=\sqrt b\sec\theta\] but this is just a given? \[x^2=\frac ba\tan^2\theta\]

OpenStudy (turingtest):

similarly we could write this as\[\sqrt{a^2x^2+b^2}\]\[x=\frac ba\tan\theta\implies x^2=\frac {b^2}{a^2}\tan^2\theta\]so we get\[\sqrt{a\left(\frac {b^2}{a^2}\tan^2\theta\right)+b^2}=\sqrt{b^2\tan^2\theta+b^2}=b\cdot\sqrt{\tan^2+1}=b\sec\theta \]I guess it looks nicer that way

OpenStudy (turingtest):

yes, it is a technique that always works; we remember this substitution as the method that will always work, as we have just seen

OpenStudy (anonymous):

ok...I'll plug in numbers of a and b and see what happens

OpenStudy (anonymous):

no way! 100 sec\theta?

OpenStudy (turingtest):

like I said, works every time :)

OpenStudy (turingtest):

wait, b for you should be 10

OpenStudy (anonymous):

\[tan^2 \theta+100= 10sec^2\theta\]

OpenStudy (turingtest):

no that statement is not true...

OpenStudy (turingtest):

what we did was say that\[\sqrt{(10\tan\theta)^2+100}=10\sec\theta\]and we did that by making a specific substitution

OpenStudy (turingtest):

in your case\[\sqrt{x^2+100}=\sqrt{x^2+10^2}\]we have\(a=1\) and \(b=10\)

OpenStudy (anonymous):

\[10tan^2 \theta+100= 10sec\theta\]

OpenStudy (turingtest):

still not quite

OpenStudy (turingtest):

can you prove that statement above?

OpenStudy (turingtest):

make a clear distinction between\[(10\tan\theta)^2=100\tan^2\theta\]and\[10\tan^2\theta\]

OpenStudy (anonymous):

\[\sqrt{(10\tan\theta)^2+100}=10\sec\theta\] \[{(10\tan\theta)^2+100}=100\sec^2\theta\]

OpenStudy (anonymous):

we're going in circles \[{100\tan^2\theta+100}=100\sec^2\theta\] \[\tan^2\theta +1=sec^2\theta\]

OpenStudy (turingtest):

this is not circles... and that is not the right simplification

OpenStudy (anonymous):

ok back to the question \[\int \frac{x^3}{\sqrt(x^2 +100)} dx\] there must some other way to solve it

OpenStudy (anonymous):

really?

OpenStudy (turingtest):

we want an expression we can integrate, and to do that we need to get rid of the square root this trig identity is how we do it, but we need to make the right sub to make it work

OpenStudy (anonymous):

Ok

OpenStudy (turingtest):

\[(10\tan\theta)^2+100=100\tan^2\theta+100=100(\tan^2\theta+1)=100\sec^2\theta\]

OpenStudy (turingtest):

now you have all that under the square root, so that is\[\sqrt{(10\tan\theta)^2+100}=\sqrt{100\tan^2\theta+100}=\sqrt{100}\cdot\sqrt{\tan^2\theta+1}=\sqrt{100\sec^2\theta}\]\[=10\sec\theta\]so hopefully you can see we have gotten somewhere

OpenStudy (turingtest):

so you have\[{x^3dx\over\sqrt{x^2+100}}\]we do the sub\[x=10\tan\theta\implies dx=10\sec^2\theta d\theta\]and we get\[{(10\tan\theta)^3{(\cancel{10}\sec^{\cancel2}\theta)}d\theta\over\cancel{10\sec\theta}}=10^3\tan^3\theta\sec\theta d\theta\]

OpenStudy (anonymous):

OH MY GOD!!!!!

OpenStudy (turingtest):

what, did it click?

OpenStudy (anonymous):

Respect

OpenStudy (anonymous):

yeah

OpenStudy (turingtest):

I'm happy respect :)

OpenStudy (anonymous):

i was trying to change the radical in a different unreasonable way

OpenStudy (turingtest):

ya know, I think this integral can be done without a trig sub, but if they really want you to do it that way...

OpenStudy (anonymous):

I don't think they need a trig substitution.... but I was trying to change \[\sqrt(tan^2\theta +100)\] to some form of \[\sqrt(tan\theta)\]....ok i guess life goes on....

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

but \[x=10\tan\theta\] instead of tan^2 theta

OpenStudy (turingtest):

actually my little trick didn't pan out, but there still may be a way to do this without a trig sub so can you handle the integral ?

OpenStudy (turingtest):

yes, but they have x^2 in the denom

OpenStudy (anonymous):

yeah the integral shouldn't be problem

OpenStudy (turingtest):

so that makes it tan^2

OpenStudy (anonymous):

thanks again Turing :)

OpenStudy (turingtest):

welcome :)

OpenStudy (anonymous):